INVESTIGATION ON REVERSE WATER-GAS SHIFT OVER La_{2}NiO_{4} CATALYST BY CW-CAVITY EN-HANCED ABSORPTION SPECTROSCOPY DURING CH_{4}/CO_{2} REFORMING

B.S. LIU, Department of Chemistry, Tianjin University, Tianjin 300072, P.R. China and Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong; LING LI, \underline{A.S.-C. CHEUNG}, Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong; C.T. AU, Department of Chemistry and Center for Surface Analysis and Research, Hong Kong Baptist University, Kowloon Tong, Hong Kong.

Reverse water-gas shift (RWGS) reaction over La_{2}NiO_{4} catalyst was investigated during CH_{4}/CO_{2} reforming by continuous wave (cw) cavity enhanced absorption spectroscopy (CEAS), a sensitive absorption technique. A cw-diode laser with wavelength in the 1.3 μm region and an optical cavity consists of a pair of high reflectivity mirrors (0.99995) were used in the experiment. Absorption lines of the ΔJ = 0 (Q branch) rotational transitions of the 2_{0}^{1}3_{0}^{1} vibrational transition of CH_{4} near 7510 cm^{−1}, and the J = 5, K_{a} = 0, K_{c} = 5 to J = 6, K_{a} = 2, K_{c} = 4 rotational transition of the 1_{0}^{1}3_{0}^{1} vibrational transition of H_{2}O near 7511.29 cm^{−1} were monitored. The results indicated that the RWGS reaction promoted the conversion of CO_{2} and decreased the partial pressure of hydrogen. Our CEAS result confirms that the conversion of CO_{2} is always 3-5% higher than CH_{4} due to RWGS reaction. In addition, the effect of reaction temperature on RWGS was studied. The work described here was supported by a grant from the Research Grants Council of Hong Kong SAR, China (Project No. HKU7101/02P)